OpenStudy (aravindg):

solve sec theta-cosec theta=4/3

OpenStudy (aravindg):

help

OpenStudy (chaise):

I wish I was more familiar with secant and cosecant, try putting this into wolframalpha

OpenStudy (anonymous):

\[Sec \theta - cosec \theta = 4/3\] correct.....

OpenStudy (aravindg):

helpppppppppp

OpenStudy (aravindg):

AEQEL HELP

OpenStudy (anonymous):

Secant and Cosecant are the inverse of the trignometric functions. So, [sin(Theta)]^-1 - [cosine(Theta)]^-1 = 4/3 Inverse both sides you get: sin(theta) - cos(theta) = 3/4 I think..

OpenStudy (aravindg):

SO???

OpenStudy (anonymous):

Do you wish to solve for theta?

OpenStudy (aravindg):

ya

OpenStudy (anonymous):

I'm sure there's a trig identity for this... Let me check my notes.

OpenStudy (aravindg):

:)

OpenStudy (aravindg):

U GUD AT trigonometry??

OpenStudy (anonymous):

lol who knows?

OpenStudy (anonymous):

I can't find my books! ><

OpenStudy (aravindg):

oops

OpenStudy (anonymous):

I did this like 2 years ago..

OpenStudy (aravindg):

thn here is another one

OpenStudy (aravindg):

ax/cos theta)+(by/sin theta)=a^2-b^2 and (axsin theta/cos^2 theta)-(bycos theta/sin^2 theta)=0 show that (ax)^2/3 +(by)^2/3=(a^2-b^2)^2/3

OpenStudy (anonymous):

apparently the correct answer is \[2\pi-\tan^{-1}(2-\sqrt{7})\]or \[2\pi-\tan^{-1}(2+\sqrt{7})\]

OpenStudy (aravindg):

how???

OpenStudy (anonymous):

these problems are for the birds

OpenStudy (aravindg):

how u got it?/?

OpenStudy (anonymous):

where do they come from?

OpenStudy (aravindg):

aeqel help plzz

OpenStudy (aravindg):

:)

OpenStudy (anonymous):

yeah i think i got it

OpenStudy (aravindg):

wow

OpenStudy (aravindg):

tl plzzzzz

OpenStudy (aravindg):

.........

OpenStudy (aravindg):

....................................

OpenStudy (anonymous):

Okay so you know that sin(theta) = cos(theta-pi/2) Replace one of the sin with that then solve.

OpenStudy (aravindg):

wt abt second question?

OpenStudy (anonymous):

btw there's another trig idenity that you need... let me find it.

OpenStudy (anonymous):

sin(2theta) = 2sin(theta)cos(theta) and cos(2theta)=cos^2(theta)-sin^2(theta)

OpenStudy (aravindg):

thn???

OpenStudy (anonymous):

Wouldn't have a clu with the other question...

OpenStudy (aravindg):

:) anothr one thn

OpenStudy (aravindg):

solve tan theta+tan 2 theta+tan 3 theta =0

OpenStudy (anonymous):

tan^2 theta ?

OpenStudy (anonymous):

or is that tan(2theta) ?

OpenStudy (aravindg):

tan(2theta)

OpenStudy (anonymous):

okay, hmmm...

OpenStudy (aravindg):

.................

OpenStudy (anonymous):

I suppose you could use the trivial solution and that is when sin(theta) = 0 and cos(theta) = 0 ..

OpenStudy (anonymous):

sorry costheta = 1

OpenStudy (aravindg):

?

OpenStudy (anonymous):

Well tan = sin/cos ... So if sin = 0, sin/cos = 0

OpenStudy (anonymous):

i.e. the Answer is theta = 0, x*pi where x is an integer

OpenStudy (anonymous):

0 or x*pi that is.

OpenStudy (aravindg):

:)

OpenStudy (anonymous):

That one is conceptual.

OpenStudy (anonymous):

i have a solution

OpenStudy (aravindg):

?

OpenStudy (anonymous):

to the first problem \[\sec(\theta)-\csc(\theta)=\frac{4}{3}\]

OpenStudy (anonymous):

actually there is something wrong with my solution, let me see if i can fix it